Operations Report – January 5th

Howdy Mission Support,

Crew 202 Operations Report 05-JAN-2019

SOL: 7

Name of person filing report: Kasey Hilton

Non-nominal systems: EVA airlock light

Notes on non-nominal systems: EVA airlock light is out, guessing the light needs to be changed

Generator (hours run): 18hr 5min; Turned on last night (04Jan2019) at 15:02; Turned off this morning (05Jan2019) at 9:07; Turned on tonight (05Jan2019) at 15:05

Solar SOC – Turned on (04Jan2019) 81%; Turned off (05Jan2019) 100%; Turned on (05Jan2019) 80%

Diesel Reading – 73%

Propane Reading – 45%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 75%; 375 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 75%; 225 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:29 to refill tank

Water Meter: 01398288 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 65.1

Beginning charge: 100%

Ending charge: 45%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 63.3

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATVs were started and run to prevent the batteries from freezing

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: The trash has not been taken out since 02Jan2019 and is starting to build up a bit in the engineering airlock

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: EVA airlock light is out, I am planning on replacing the bulb in the morning; No change in power system issues from earlier this week, still having to manually start the generator and the SOC is not increasing

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Best,

Kasey